OpenStudy (anonymous):

If the length, width, and height of a rectangular prism is multiplied by 5 will it be *5, *15, *25, *125

OpenStudy (anonymous):

How will the volume change? **

OpenStudy (anonymous):

Do you know how to find the volume of a prism?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

.... sorry... do you know the formula for the volume of a prism?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Volume = ???? what's the formula?

OpenStudy (anonymous):

L*W*H is formula.

OpenStudy (anonymous):

ok... so Volume = L * W * H now since each dimension is multiplied by 5, we have Volume = (5*L) * (5*W) * (5*H) can you simplify that form me?

OpenStudy (anonymous):

So it's 5 times greater?

OpenStudy (anonymous):

no.... how did you get that?

OpenStudy (anonymous):

5*L 5*W 5*H It doesn't give us l,w, and h in the question.

OpenStudy (anonymous):

but if you simplify it you'll see how the volume changed.....

OpenStudy (anonymous):

It doesn't give us L,W, and H.

OpenStudy (anonymous):

yes... L, W, H can be any number, but that's ok... simplify: New Volume = 5L*5W*5H

OpenStudy (anonymous):

multiply the numbers....

OpenStudy (anonymous):

25 times greater. Thank you. :D

OpenStudy (anonymous):

New Volume = 5L*5W*5H = 5*5*5*L*W*H = 125 * L*W*H = 125 * (original volume)

OpenStudy (anonymous):

.. 62^3 original 1550^3 new 62 * 25 = 1550

OpenStudy (anonymous):

125 not 25...

OpenStudy (anonymous):

Well that's what I got so yeah.

OpenStudy (anonymous):

ok...

Directrix (directrix):

@AyeItsEmily The original rectangular prism will be similar to the one formed by multiplying the original prism's l, w, and h by 5. Recall the theorem: The ratio of volumes of two similar figures is the cube of the scale factor. The scale factor from the original solid to the enlarged similar one is 1 to 5 or 1/5. So, (1/5) ^ 3 = ratio of volumes 1/125 is the ratio of the volumes. So, the volume of the original rectangular prism is multiplied by a factor of 125 to get the volume of the enlarged prism.